OpenStudy (anonymous):

4x^3-8x = 0 solve for x

OpenStudy (kagıtucak):

4x(x^2-2)=0 4x=0 x=0 or x^2-2=0 x^2=2 x=+(2^1/2) or x=-(2^1/2)

OpenStudy (anonymous):

4x^3-8X=0 X(4X^2-8)=0 4X^2-8=0 DIVIDED BY X 4X^2-8=0 4(X^2-2)=0 X^2-2=0 DIVIDED BY 4 X^2-2=0 X^2=2 X=SQUARE ROOT OF 2